# Microscopic Structure and Dynamics of Interfacial Water at Fluorinated vs Nonfluorinated Surfaces - Insights from Ab-Initio Simulations and IR Spectroscopy

### 27 Mar 2026 v1

@@ II.2 Linear IR spectroscopy of interfacial water
The position and line shape of the OH stretch peak of water provide valuable insights into the hydrogen bond network of water, as both parameters are known to depend on hydrogen bond strength Nakamoto, Margoshes, and Rundle (1955). It has been argued that in liquid water, the OH stretch peak can be decomposed into contributions from populations with different numbers of hydrogen bonds Bergonzi et al. (2014), allowing for the tracking of the hydrogen bonding structure. While the uniqueness of such a decomposition based on fitting experimental spectra is questionable due to the large spectral overlap of different contributions, but also the intricate nature of hydrogen bonding environments Auer et al. (2007); Tainter et al. (2013), the overall interpretation that a higher frequency of the OH stretch vibration is related to a lower degree of hydrogen-bonding and a lower frequency to a higher degree of hydrogen-bonding is largely accepted Calegari Andrade and Pham (2023); Wang et al. (2025). Here, we determine the center frequency of the OH stretch band by fitting to a single Gaussian function
@@ II.3 Free OH groups as spectroscopic fingerprints of water-surface interactions
To experimentally test our predictions regarding the frequency of free OH vibrations, we conduct surface enhanced infrared absorption spectroscopy experiments at the HSAM-water and FSAM-water interfaces. As model molecules, we used 1-hexanethiol and 1,1,2,2-tetrahydrogen-perfluorohexane-1-thiol, both with a molecular length of six carbon atoms (see Fig. 4 insets). The interfacial water structure was measured by depositing a monolayer of the HSAM or FSAM compound directly from water. To increase the solubility of these compounds, we used an alkaline buffer (pH = 9.5) and kept the concentration low during deposition (1.5 mM). First, a background spectrum was recorded by covering the bare gold surface with the alkaline buffer. The surface enhancement factor characteristic of SEIRAS is restricted to a distance of about 100 nm from the gold surface Tseng et al. (2023). This allows us to observe the vibrational response of molecules within this distance and removes the strong contribution from the bulk water above the monolayer. The electric field in SEIRAS is polarized perpendicular to the gold surface, enhancing vibrational modes with transition dipole moments oriented in this direction. This makes SEIRAS particularly suitable for detecting signatures of free OH groups Ataka, Yotsuyanagi, and Osawa (1996), which were observed only in the perpendicular IR spectra (Fig. 3 D,F). After adding the thiol solution, the gold substrate gets quickly covered with the HSAM/FSAM molecules. As the SAM forms, water molecules are displaced away from the gold interface. This removes the spectroscopic signature from the gold-water interface and the contribution from the liquid water volume occupied by the monolayer. After 20 s, surface coverage reaches saturation, which is observed by a stabilization in the intensity of the bands at 2950 cm-1 and the bands at 1250 cm-1 of the HSAM and FSAM cru (2025), respectively (Fig. 4). In both experiments, the broad O-H stretching vibrational band of water is observed at 3285 cm-1. The fact that the band appears negative relates to the displacement of liquid bulk water by the monolayer volume and from the lost water structure at the gold interface. At higher frequencies, a sharp positive band is observed at 3682 cm-1 for the FSAM and at 3672 cm-1 for the HSAM, characteristic frequencies of free O-H groups of the newly created water structure at the water-monolayer interface. This dangling-water band appears notably sharper in the FSAM. These results confirm our simulation predictions: free OH vibrations are blue shifted at the FSAM-water interface with respect to the HSAM-water interface.
@@ II.4 The line shape of free OH spectra reveal a slowdown of water interfacial dynamics at fluorinated SAMs
Results are shown in Fig. 7 E-G. Indeed, the lifetime of free OH groups is longer at the FSAM-water interface than at the air-water and HSAM-water interfaces, as indicated by the slower decay of the correlation function in Fig. 7 E. Since our criterion for free OH groups involves both and , comparing (Fig. 7 E) with (Fig. 7 F) reveals that free OH lifetimes are determined approximately equally by translational motion reorientation. Moreover, not only are free OH dynamics slowed at the FSAM–water interface relative to the other two, but the translational motion of hydration water in general is also decelerated, as shown by in Fig. 7 G. While the interfacial water at strongly hydrophobic surfaces exhibits slower reorientational dynamics than that at surfaces of intermediate hydrophobicity, similar to the observed slowdown of dynamics around hydrophobic solutes Rezus and Bakker (2007), very strong slowdowns are typically observed in the vicinity of strongly polar surfaces and thus associated with hydrophilic interfaces Romero-Vargas Castrillón et al. (2009); Stirnemann et al. (2011); Gekle and Netz (2012). This leads to the intriguing conclusion that fluorinated SAMs, despite being macroscopically more hydrophobic than their unfluorinated counterparts, exhibit microscopic characteristics usually associated with increased hydrophilicity.
@@ III Conclusion
Finally, we demonstrate that the line shape of the free OH spectral peak conveys information about the dynamics of water in the immediate vicinity of the interface. The line width of these features is determined by the lifetime of free OH groups, which, in turn, is similarly limited by the rotational and translational motion of water molecules. Interestingly, the lifetime of free OH groups is considerably longer at the FSAM-water interface than at the air-water and HSAM-water interfaces, resulting in a narrow free OH peak. This lifetime is also correlated with an overall slowing down of water dynamics at the FSAM-water interface. While some degree of deceleration of water dynamics is typical for hydrophobic interfaces Romero-Vargas Castrillón et al. (2009), a slowdown of water dynamics as strong as the one observed in our simulations is typically associated with hydrophilic interfaces, rendering fluorinated SAMs rather atypical hydrophobic surfaces.
@@ IV Methods
DFT-MD simulations of the SAM-water systems are carried out with the CP2K/7.1 software suite Kühne et al. (2020) using the BLYP exchange-correlation functional Becke (1988); Lee, Yang, and Parr (1988). The electronic valence density is expanded in the MOLOPT-SR-DZVP basis set VandeVondele and Hutter (2007), which is treated in reciprocal space using a plane-wave cutoff of 400 Ry. Atomic cores including core electrons are modeled using Goedecker-Teter-Hutter pseudopotentials Goedecker, Teter, and Hutter (1996). MD simulations are run using a time step of 0.5 fs for at least 100 ps each, after an additional equilibration time of 10 ps.
The simulations are conducted in the NpT ensemble, employing the canonical sampling through velocity rescaling algorithm Bussi, Donadio, and Parrinello (2007) to maintain the systems at a constant temperature of 300 K. To maintain a constant pressure in the system, we implement the following piston setup: Carbon atoms at the bottom of the SAMs are restrained to a hexagonal grid (Fig. 1 D-E) using harmonic restraint potentials of the form
@@
A universal single-reflection silicon attenuated total-reflection (ATR) crystal from IRubis was covered with a SEIRAS-active gold layer using the chemical deposition method from Miyake et al. Miyake, Ye, and Osawa (2002). The silicon surface was first treated with a 1:1:1 (v/v/v) hot solution (90 ) of HCl (37%, Carl Roth), (30%, Chemsolute), and water. The surface was then polished with the help of powder (50 nm, Sigma Aldrich) and treated with a (40%, Fluka) solution for one minute. Gold deposition was performed by covering the surface with a 1:1:1 (v/v/v) mixture of:

### 03 Apr 2026 v2

@@ II.2 Linear IR spectroscopy of interfacial water
The position and line shape of the OH stretch peak of water provide valuable insights into the hydrogen bond network of water, as both parameters are known to depend on hydrogen bond strength Nakamoto et al. (1955). It has been argued that in liquid water, the OH stretch peak can be decomposed into contributions from populations with different numbers of hydrogen bonds Bergonzi et al. (2014), allowing for the tracking of the hydrogen bonding structure. While the uniqueness of such a decomposition based on fitting experimental spectra is questionable due to the large spectral overlap of different contributions, but also the intricate nature of hydrogen bonding environments Auer et al. (2007); Tainter et al. (2013), the overall interpretation that a higher frequency of the OH stretch vibration is related to a lower degree of hydrogen-bonding and a lower frequency to a higher degree of hydrogen-bonding is largely accepted Calegari Andrade and Pham (2023); Wang et al. (2025). Here, we determine the center frequency of the OH stretch band by fitting to a single Gaussian function
@@ II.3 Free OH groups as spectroscopic fingerprints of water-surface interactions
To experimentally test our predictions regarding the frequency of free OH vibrations, we conduct surface enhanced infrared absorption spectroscopy experiments at the HSAM-water and FSAM-water interfaces. As model molecules, we used 1-hexanethiol and 1,1,2,2-tetrahydrogen-perfluorohexane-1-thiol, both with a molecular length of six carbon atoms (see Fig. 4 insets). The interfacial water structure was measured by depositing a monolayer of the HSAM or FSAM compound directly from water. To increase the solubility of these compounds, we used an alkaline buffer (pH = 9.5) and kept the concentration low during deposition (1.5 mM). First, a background spectrum was recorded by covering the bare gold surface with the alkaline buffer. The surface enhancement factor characteristic of SEIRAS is restricted to a distance of about 100 nm from the gold surface Tseng et al. (2023). This allows us to observe the vibrational response of molecules within this distance and removes the strong contribution from the bulk water above the monolayer. The electric field in SEIRAS is polarized perpendicular to the gold surface, enhancing vibrational modes with transition dipole moments oriented in this direction. This makes SEIRAS particularly suitable for detecting signatures of free OH groups Ataka et al. (1996), which were observed only in the perpendicular IR spectra (Fig. 3 D,F). After adding the thiol solution, the gold substrate gets quickly covered with the HSAM/FSAM molecules. As the SAM forms, water molecules are displaced away from the gold interface. This removes the spectroscopic signature from the gold-water interface and the contribution from the liquid water volume occupied by the monolayer. After 20 s, surface coverage reaches saturation, which is observed by a stabilization in the intensity of the bands at 2950 cm-1 and the bands at 1250 cm-1 of the HSAM and FSAM Cruz et al. (2025), respectively (Fig. 4). In both experiments, the broad O-H stretching vibrational band of water is observed at 3285 cm-1. The fact that the band appears negative relates to the displacement of liquid bulk water by the monolayer volume and from the lost water structure at the gold interface. At higher frequencies, a sharp positive band is observed at 3682 cm-1 for the FSAM and at 3672 cm-1 for the HSAM, characteristic frequencies of free O-H groups of the newly created water structure at the water-monolayer interface. This dangling-water band appears notably sharper in the FSAM. These results confirm our simulation predictions: free OH vibrations are blue shifted at the FSAM-water interface with respect to the HSAM-water interface.
@@ II.4 The line shape of free OH spectra reveal a slowdown of water interfacial dynamics at fluorinated SAMs
Results are shown in Fig. 7 E-G. Indeed, the lifetime of free OH groups is longer at the FSAM-water interface than at the air-water and HSAM-water interfaces, as indicated by the slower decay of the correlation function in Fig. 7 E. Since our criterion for free OH groups involves both and , comparing (Fig. 7 E) with (Fig. 7 F) reveals that free OH lifetimes are determined approximately equally by translational motion reorientation. Moreover, not only are free OH dynamics slowed at the FSAM–water interface relative to the other two, but the translational motion of hydration water in general is also decelerated, as shown by in Fig. 7 G.
Interfacial water at strongly hydrophobic surfaces is known to exhibit slower reorientational dynamics than at surfaces of intermediate hydrophobicity, primarily due to enhanced water–water interactions, similar to the slowdown observed around hydrophobic solutes Rezus and Bakker (2007); Romero-Vargas Castrillón et al. (2009); Stirnemann et al. (2011). Since the air–water interface represents the most hydrophobic limit, the pronounced slowdown observed at HSAM and FSAM surfaces must involve an additional mechanism, most likely a retardation arising from surface–water interactions. Such a process is typically associated with strongly polar, and thus hydrophilic, interfaces, whose interpretation is also more consistent with the magnitude of the observed slowdown Stirnemann et al. (2011); Gekle and Netz (2012). This leads to the intriguing conclusion that fluorinated SAMs, despite being macroscopically more hydrophobic than their unfluorinated counterparts, exhibit also some spectroscopic characteristics usually associated with increased hydrophilicity.
@@ III Conclusion
Finally, we demonstrate that the line shape of the free OH spectral peak conveys information about the dynamics of water in the immediate vicinity of the interface. The line width of these features is determined by the lifetime of free OH groups, which, in turn, is similarly limited by the rotational and translational motion of water molecules. Interestingly, the lifetime of free OH groups is considerably longer at the FSAM-water interface than at the air-water and HSAM-water interfaces, resulting in a narrow free OH peak. This lifetime is also correlated with an overall slowing down of water dynamics at the FSAM-water interface. While some degree of dynamical deceleration is typical for hydrophobic interfaces Romero-Vargas Castrillón et al. (2009), the pronounced slowdown observed in our simulations is rather associated with hydrophilic interfaces, rendering fluorinated SAMs atypical hydrophobic surfaces from a spectroscopic perspective.
@@ IV Methods
DFT-MD simulations of the SAM-water systems are carried out with the CP2K/7.1 software suite Kühne et al. (2020) using the BLYP exchange-correlation functional Becke (1988); Lee et al. (1988). The electronic valence density is expanded in the MOLOPT-SR-DZVP basis set VandeVondele and Hutter (2007), which is treated in reciprocal space using a plane-wave cutoff of 400 Ry. Atomic cores including core electrons are modeled using Goedecker-Teter-Hutter pseudopotentials Goedecker et al. (1996). MD simulations are run using a time step of 0.5 fs for at least 100 ps each, after an additional equilibration time of 10 ps.
The simulations are conducted in the NpT ensemble, employing the canonical sampling through velocity rescaling algorithm Bussi et al. (2007) to maintain the systems at a constant temperature of 300 K. To maintain a constant pressure in the system, we implement the following piston setup: Carbon atoms at the bottom of the SAMs are restrained to a hexagonal grid (Fig. 1 D-E) using harmonic restraint potentials of the form
@@
A universal single-reflection silicon attenuated total-reflection (ATR) crystal from IRubis was covered with a SEIRAS-active gold layer using the chemical deposition method from Miyake et al. Miyake et al. (2002). The silicon surface was first treated with a 1:1:1 (v/v/v) hot solution (90 ) of HCl (37%, Carl Roth), (30%, Chemsolute), and water. The surface was then polished with the help of powder (50 nm, Sigma Aldrich) and treated with a (40%, Fluka) solution for one minute. Gold deposition was performed by covering the surface with a 1:1:1 (v/v/v) mixture of: